OpenStudy (anonymous):

How can 2x2 = x2 + 4x + 8 be set up as a system of equations?

OpenStudy (whpalmer4):

Quite easily! It is identical in effect to the system\[y = 2x^2\]\[y = x^2+4x+8\]Any solution to that system of is going to have equal values of \(y\) for equal values of \(x\), right?

OpenStudy (whpalmer4):

and if we substitute using either equation: \[y = 2x^2\]substitute in \(x^2+4x+8\) for \(y\) \[x^2+4x+8 = 2x^2\]That's the same as what you have, and likewise \[y = x^2+4x+8\]substitute in \(y = 2x^2\)\[2x^2=x^2+4x+8\]